OpenStudy (anonymous):

Not understanding how to work sum of cubes 125x^3+27y^3

OpenStudy (anonymous):

do u need to factor them

OpenStudy (anonymous):

then use a^3+b^3= (a+b)(a^2 -ab+b^2)

OpenStudy (anonymous):

for a^3-b^3= (a-b)(a^2 +ab+b^2)

OpenStudy (agent0smith):

\[ 125x^3+27y^3 = (5x)^3 + (3y)^3\]

OpenStudy (anonymous):

if u need not factor them then use a^3+b^3 =(a+b)^3 -3ab(a+b)

OpenStudy (agent0smith):

Now a = 5x and b =3y.

OpenStudy (anonymous):

125x^3+27y^3 =(5x)^3 +(3y)^3 =(5x+3y)(25x^2 -15xy +9y^2)